Determine the Resultant Vector of the following vectors:
A = 26.5 m 56° NW; B = 44.0 m 28° NE; and C = 31.0 m S.
1
Expert's answer
2020-10-21T06:29:20-0400

The coordinates of the vectors are:


"\mathbf{A} = 26.5\times (-\sin56\degree,\cos56\degree) \approx (-21.97, 14.82)\\
\mathbf{B} = 44\times (\sin28\degree,\cos28\degree)\approx (20.66,38.85)\\
\mathbf{C} = 31\times (0,-1) = (0,-31)"

Their sum is:


"\mathbf{A} +\mathbf{B}+\mathbf{C} = \\
=(-21.97+20.66+0,\space14.82+38.85-31)= (-1.31, 22.67)"

Answer. "(-1.31, 22.67)".